OpenStudy (anonymous):

what is the next item in the sequence: 1,-2,-5,-8,-11?

OpenStudy (anonymous):

Its -14 right?

OpenStudy (anonymous):

Common difference is -3 So every term has a gap of decreasing 3 Yes you are right..

OpenStudy (helder_edwin):

it looks like an arithmetic progression \[ -2-1=-3 \] \[ -5-(-2)=-3 \] \[ -8-(-5)=-3 \] \[ -11-(-8)=-3 \]

OpenStudy (helder_edwin):

yes u r right